# High-Resolution Flood Probability Mapping Using Generative Machine Learning with Large-Scale Synthetic Precipitation and Inundation Data

### v1

@@ Abstract
High-resolution flood probability maps are essential for addressing the limitations of existing flood risk assessment approaches but are often limited by the availability of historical event data. Also, producing simulated data needed for creating probabilistic flood maps using physics-based models involves significant computation and time effort inhibiting the feasibility. To address this gap, this study introduces Flood-Precip GAN (Flood-Precipitation Generative Adversarial Network), a novel methodology that leverages generative machine learning to simulate large-scale synthetic inundation data to produce probabilistic flood maps. With a focus on Harris County, Texas, Flood-Precip GAN begins with training a cell-wise depth estimator using a limited number of physics-based model-generated precipitation-flood events. This model, which emphasizes precipitation-based features, outperforms universal models. Subsequently, a Generative Adversarial Network (GAN) with constraints is employed to conditionally generate synthetic precipitation records. Strategic thresholds are established to filter these records, ensuring close alignment with true precipitation patterns. For each cell, synthetic events are smoothed using a K-nearest neighbors algorithm and processed through the depth estimator to derive synthetic depth distributions. By iterating this procedure and after generating 10,000 synthetic precipitation-flood events, we construct flood probability maps in various formats, considering different inundation depths. Validation through similarity and correlation metrics confirms the fidelity of the synthetic depth distributions relative to true data. Flood-Precip GAN provides a scalable solution for generating synthetic flood depth data needed to create high-resolution flood probability maps, significantly enhancing flood preparedness and mitigation efforts.
Lipai Huanga, *, Federico Antolinib, Ali Mostafavia,b, Russell Blessingb, Matthew Garciac, Samuel D. Brodyb
a Urban Resilience.AI Lab, Zachry Department of Civil and Environmental Engineering,
Texas A&M University, College Station, TX
b Institute for a Disaster Resilient Texas,
Texas A&M University, College Station, TX
c Civil and Environmental Engineering,
Rice University, Houston, TX
* corresponding author, email: lipai.huang@tamu.edu
Keywords Flood Risk $\cdot$ Generative Modeling $\cdot$ Machine Learning $\cdot$ Synthetic Data $\cdot$ Data Augmentation $\cdot$ Precipitation Events Generation
@@ 1 Introduction
Flood hazards pose a significant threat to cities and communities globally, resulting in extensive physical damage and substantial economic costs due to material losses and human casualties, particularly in densely populated regions. In the United States, flood-related insurance claims have consistently averaged around \$40 billion annually over the past four decades [1]. With the projected increase in the frequency and severity of flood events across various U.S. regions [2, 3], this financial burden is expected not only to persist but also to escalate. The prompt and accurate assessment of flood risk is crucial for the efficient response and recovery of affected communities. Spatial flood risk characterization traditionally involves the creation of floodplain maps, which delineate areas typically prone to flooding based on historical data and hydrological models. These maps are crucial for flood risk reduction, urban planning, and development regulations, as they identify regions where flood risk management strategies are most needed. However, floodplain maps often possess limitations due to their reliance on limited historical data, which may not accurately reflect current conditions. Also, development of floodplains using physics-based hydraulic and hydrologic (H&H) models is computationally expensive. In addition, these maps usually depict flood extent limited to a specific expected frequency of flooding, limiting their utility in risk assessments and communication. The gaps in traditional floodplain mapping underscore the need to incorporate probability into flood maps beyond the traditional 100-year and 500-year frequencies. In the context of this study, probabilistic flood maps quantify flood risk irrespective of the likelihood of rainfall events. That is, probabilistic flood maps estimate the likelihood that a rainfall event causes a certain inundation in a given place (e.g., 1 ft) from a large number of rainfall events, without considering the probability of occurrence of each of those. For example, if 350 out of 1000 rainfall events cause at least 1 ft inundation in a certain place, a probabilistic flood map will associate a likelihood of 0.35 to the 1-ft inundation of that place for rainfall events with the range of characteristics considered. This quantified likelihood provides a measure of 1-ft inundation risk in the area given the characteristics of rainfall events. In this context, this probability is different from the annual likelihood typically used which is based on the return period of rainfall events. By offering a probabilistic measure of flood risks (based on characteristics of rainfall events but independent of the return period of rainfall events), maps provide quantitative insights that help policymakers, planners, and communities to better prepare for and mitigate the impacts of potential flooding events with tailored, data-driven strategies. The creation of flood probability maps hinges on simulating a large number of flood inundation scenarios based on which the likelihood of a range of flood depths across a region can be estimated. In fact, one of the major barriers to creation of flood probability maps is the lack of numerous flood scenarios. On one hand, the limited number of historical flood events in a given region is not large enough to build a robust statistical distribution of rainfall events and their associated flood depths. Also, depending on the rate of land use change, historical data many not be adequate to estimate flood risk in current conditions. On the other hand, generating a large number of simulated flood scenarios using H&H models is computationally expensive and would not be feasible. An alternative to historical data is represented by synthetic high-resolution flood scenarios generated using generative models.
@@
The use of generative models and machine learning has gained significant attention in flood risk analysis. Recent studies have adopted Generative Adversarial Networks (GANs) to improve data generation, such as [4], which uses GANs to generate synthetic precipitation data integrated with the Soil and Water Assessment Tool (SWAT) to enhance flood frequency analysis. Similarly, [5] employs machine learning models within a GAN framework for real-time flood forecasting in urban environments, improving prediction accuracy and emergency response strategies through event-based data modeling. Deep learning models, including Convolutional Neural Networks (CNNs) and Recurrent Neural Networks (RNNs), have been used for predicting flash flood probabilities and integrating with hydrodynamic models for efficient flood mapping and risk assessment [6]. Machine learning algorithms like Random Forest and Extreme Gradient Boosting (XGBoost) have been applied to assess flood susceptibility and generate flood risk maps by analyzing key flood conditioning factors [7, 8]. Furthermore, studies such as [9, 10] have developed ML-based models, including Long Short-Term Memory (LSTM) networks and conditional GANs, to enhance flood prediction accuracy in specific urban settings. Despite these advancements, there is currently no model for simulating synthetic flood data specifically for creating probabilistic flood maps. To address these challenges, we introduce Flood-Precip GAN, a novel methodology leveraging generative machine learning to synthesize large-scale inundation data. The main workflow is illustrated in Fig. 1. Flood-Precip GAN starts with the training of a depth estimator using a number of physics-based model-generated flood events, incorporating precipitation-based, spatial, and region-specific features. Following this, a Conditional GAN (CTGAN) with constraints is employed to generate synthetic rainfall precipitation events. Strategic thresholds are established to filter these synthetic records, ensuring close alignment with true precipitation patterns. For each location, synthetic events are smoothed using a K-nearest neighbors (KNN) algorithm and processed through the depth estimator to derive synthetic depth distributions. By iterating this procedure across multiple synthetic rainfall events, we construct flood probability maps for different rainfall categories with certain duration, maximum, and cumulative precipitation ranges. These maps are validated using similarity and correlation metrics to confirm the correspondence of synthetic depth distributions to true data. The novel contributions of the model presented in this paper are threefold:
@@
    We introduce a novel methodology that utilizes a surrogate machine learning pipeline to estimate flood probabilities using synthetic precipitation-flood events generated by CTGAN.
@@
Figure 1: The main research workflow consists of four steps. Step I involves training and selecting the optimal model for depth estimation to generate synthetic depth data. Step II focuses on training a CTGAN with specific constraints and using conditional sampling to create a synthetic point cloud. In Step III, a cell-level LMH pool is developed based on thresholds derived from sampled training events, allowing for strategic sampling of features to produce synthetic precipitation-flood events with the depth estimator. Finally, Step IV repeats III to build a synthetic depth distribution, culminating in the creation of a synthetic flood probability map.
Flood-Precip GAN provides a scalable solution for generating high-resolution flood probability maps, significantly enhancing flood preparedness and mitigation efforts. The following section explains the components of the Flood-Precip GAN model and the associated datasets.
@@ 2.1.1 Study Region
Figure 2: Study area and Harris County Flood Control District gauge distribution. Map created with ArcGIS Pro 3.0.0 (https://pro.arcgis.com/).
As Fig. 2 depicts, the test bed for creating the Flood-Precip GAN model is Harris County in Texas. The main watersheds are Cypress Creek and Buffalo Bayou. Harris County, the core county of the Greater Houston Metropolitan Statistical Area, spans 1,778 square miles (4,605 square kilometers) and has grown to over 4.5 million residents in the past decade. The county’s flat topography, with elevations ranging from -40 ft (-12.19 meters) to 300 ft (91.44 meters), and predominantly developed land (over two-thirds), along with 20% pasture and cultivated lands [11], contribute to its flood risk. The Cypress Creek watershed in the north and the Buffalo Bayou system in the central and southern parts flow into the San Jacinto River and the Ship Channel, eventually reaching the Gulf of Mexico. The area’s dense development, poor natural drainage, limited soil infiltration, and subtropical climate significantly heighten its susceptibility to chronic flooding.
@@ 2.1.2 Data
In creating Flood-Precip GAN, there is a need for data related to flood inundation scenarios from a range of rainfall events. The lack of sufficient historical inundation data was addressed using 592 precipitation-flood events simulated by the HEC-RAS 2D model, a hydrologic software developed by the United States Army Corps of Engineers. This method, leveraging the advantages of the Rasterized Time-series Resampling Method (RTRM), is a non-denationalized bootstrapping approach for Radar Rainfall data that enhances input space coverage, captures spatiotemporal patterns, supports both temporal and spatial scaling, maintains the integrity of input dataset patterns, and facilitates interpolation to fill data gaps, providing smoother and more accurate coverage for hydrological modeling. The study area was modeled using a 1,200 ft $\times$ 1,200 ft (approximately 366 m) mesh grid, which was further refined along major watercourses and tributaries to improve computational stability and accuracy. The resulting mesh comprised 26,301 cells, providing a detailed representation of the area. Parameters such as Manning’s roughness and imperviousness were derived from the 2019 National Land Cover Dataset [11], while soil infiltration values were sourced from the Gridded Soil Survey Geographic database [12]. Model validation involved analyzing data from Harris County Flood Control District gauges [13] during five significant flood events from 2016 to 2020, including the 2016 Tax Day flood and Hurricane Harvey in 2017. This validation process, which calibrated the HEC-RAS 2D model for global minimum error across all gauges, confirmed the model’s reliability and accuracy for simulating diverse flood scenarios [14]. The rainfall events used in generating flood scenarios by HEC-RAS 2D are further used to generate synthetic rainfall events as described in the following sections. The generated flood depth scenarios were used to create the depth estimator (a surrogate machine learning model for determining maximum flood depth) given rainfall and other landscape features.
@@ 2.2 Preprocessing
We employed Voronoi tessellation [15] to partition the training dataset, assigning each point within the study area a unique polygonal region to ensure there is no overlap between cells. Additionally, we segmented the 592 rainfall events used in the HEC-RAS 2D flood scenarios into three inundation depth levels to ensure consistent stratification for subsequent dataset splitting and resampling. Specifically, for the training set, events are distributed with mean depths of $\leq 2\ \text{inches}$, between $>2$ and $\leq 6\ \text{inches}$, and $>6\ \text{inches}$ in a ratio of 2:4:1, respectively. The validation set consists of 20% of the training set volume, with the remaining data designated as the test set. This approach optimizes the separation of events to ensure relatively balanced representation across the training, validation, and testing phases, facilitating more effective modeling.
@@ 2.3 Depth Estimator Model
For the depth estimator component of Flood-Precip GAN, we chose a tree-based machine learning model trained on simulated scenarios from the HEC-RAS 2D model. Compared to traditional machine learning models, tree-based models have shown superior performance due to their ability to manage complex, non-linear relationships and interactions within the data. XGBoost [16, 17], in particular, excels with its gradient boosting framework, which iteratively improves model accuracy by minimizing residual errors from previous iterations. Additionally, deep learning models in regression, such as the Transformer [18, 19], utilize attention mechanisms to capture long-range dependencies and complex feature interactions, providing substantial benefits in predictive performance. In this study, we employed the XGBoost regressor and the regression transformer as universal depth estimators, training them on physics-based model-generated events that include both spatial features incorporating channel presence and terrain elevation, and precipitation-based features as shown in Table 1. This integration aims to enhance the accuracy and robustness of maximum flood inundation estimations. Furthermore, we compared these universal models with a cell-wise model, MaxFloodCast V2 [20], an XGBoost-based depth estimator trained specifically on individual cell meshes, resulting in 26,301 models corresponding to the 26,301 cells. Due to the unique nature of cell-wise model implementation, we assumed geospatial changes to be negligible across all events, thereby excluding spatial features. To account for the influence of inter-watershed precipitation without directly utilizing static geospatial features, we introduced cumulative and peak precipitation ratios as shown in Eq. 1 and Eq. 2 to effectively capture interactions among watersheds.
[…]
$h_{i,j}=\begin{cases}1,&\text{if }p_{i,j}>2\text{ in},\\ 0,&\text{otherwise}.\end{cases}$ (2)
where $AW_{i}$ represents the area of watershed $i$, $AC_{i,j}$ represents the area of cell $j$ in watershed $i$, $h_{i,j}$ is a binary identifier of cell $j$ in watershed $i$, and $p_{i,j}$ denotes the peak (cumulative) precipitation in cell $j$ within watershed $i$. Heavy peak precipitation ratio and heavy cumulative precipitation ratio are defined at the watershed level and thus are constant across all cells of a watershed for a given rainfall event. These two ratios can vary significantly across different events and watershed regions, providing substantial explanatory potential.
@@
9 Heavy Cumulative Precipitation Ratio (one for each watershed) Cell-wise Scalar (range 0 to 1) Proportion of watershed area experiencing heavy cumulative precipitation
9 Heavy Peak Precipitation Ratio (one for each watershed) Cell-wise Scalar (range 0 to 1) Proportion of watershed area experiencing heavy peak precipitation
We assessed the predictive performance of the depth estimators (in Section 3.1) using Root Mean Squared Error (RMSE) and R-squared ($R^{2}$) scores [21] to evaluate both accuracy and explainability. RMSE measures the average magnitude of the prediction errors, providing insight into the estimator’s precision, while $R^{2}$ indicates the proportion of variance in the observed data that is predictable from the input features, reflecting the model’s explanatory power. In addition, we conducted a comprehensive analysis of the aggregated performance of these models on both channel and non-channel cells. This distinction is crucial due to the varying hydrological and ground texture characteristics inherent to these environments, which influence the uncertainty and behavior of flood dynamics.

### v2

@@ Abstract
[Abstract]High-resolution flood probability maps are instrumental for assessing flood risk but are often limited by the availability of historical data. Additionally, producing simulated data needed for creating probabilistic flood maps using physics-based models involves significant computation and time effort, which inhibit its feasibility. To address this gap, this study introduces Precipitation-Flood Depth Generative Pipeline, a novel methodology that leverages generative machine learning to generate large-scale synthetic inundation data to produce probabilistic flood maps. With a focus on Harris County, Texas, Precipitation-Flood Depth Generative Pipeline begins with training a cell-wise depth estimator using a number of precipitation-flood events model with a physics-based model. This cell-wise depth estimator, which emphasizes precipitation-based features, outperforms universal models. Subsequently, the Conditional Generative Adversarial Network (CTGAN) is used to conditionally generate synthetic precipitation point cloud, which are filtered using strategic thresholds to align with realistic precipitation patterns. Hence, a precipitation feature pool is constructed for each cell, enabling strategic sampling and the generation of synthetic precipitation events. After generating 10,000 synthetic events, flood probability maps are created for various inundation depths. Validation using similarity and correlation metrics confirms the accuracy of the synthetic depth distributions. The Precipitation-Flood Depth Generative Pipeline provides a scalable solution to generate synthetic flood depth data needed for high-resolution flood probability maps, which can enhance flood mitigation planning.
keywords:
Flood Risk, Generative Modeling, Machine Learning, Synthetic Data, Data Augmentation, Precipitation Events Generation
articletype: ORIGINAL ARTICLE
@@ 1 INTRODUCTION
Flood hazards pose a significant threat to cities and communities globally, resulting in extensive physical damage and substantial economic costs due to material losses and human casualties, particularly in densely populated regions. In the United States, flood-related insurance claims have consistently averaged around \$40 billion annually over the past four decades 33. With the projected increase in the frequency and severity of flood events across various U.S. regions 19, 26, this financial burden is expected not only to persist but also to escalate. An accurate assessment of flood risk is crucial to plan for efficient response and recovery of potentially affected communities. Flood risk characterization traditionally involves the creation of floodplain maps, which delineate areas typically prone to flooding based on historical data and hydrologic simulations. However, floodplain maps often have limitations due to their reliance on limited historical data, which may not accurately reflect current and future conditions. Also, development of floodplains using physics-based hydraulic and hydrologic (H&H) models is computationally expensive. In addition, these maps usually depict flood extent for a few specific flooding frequency, limiting their utility in risk assessments and communication. The gaps in traditional floodplain mapping underscore the need to incorporate probability into flood maps beyond the traditional 100-year and 500-year frequencies. In the context of this study, probabilistic flood maps quantify flood risk from a distribution of rainfall events, without explicitly considering the frequency of each of those events. That is, probabilistic flood maps estimate the likelihood that a rainfall event causes a certain inundation in a given place (e.g., 1 ft) from a large number of rainfall events, without considering the probability of occurrence of each event. For example, if 350 out of 1000 rainfall events cause at least 1 ft inundation in a certain place, a probabilistic flood map will associate a likelihood of 0.35 to the 1-ft inundation of that place for rainfall events with the range of characteristics considered. This quantified likelihood provides a measure of 1-ft inundation risk in the area given the characteristics of rainfall events. In this context, this probability is different from the annual likelihood typically used, which is based on a statistical analysis of historical storms and the extrapolation of flood depth and/or extent values for a few return periods (e.g., 10-, 100-, 500-year storm). By offering a spatially continuous probabilistic measure of flood risks, maps provide quantitative insights that can help policymakers, planners, and communities in the mitigation and preparedness phases with tailored, data-driven strategies that lessen the impacts of potential flooding 1, 20.
The creation of flood probability maps hinges on simulating a large number of flood inundation scenarios based on which the likelihood of a range of flood depths across a region can be estimated. In fact, one of the major barriers to creation of flood probability maps is the lack of numerous flood scenarios. On one hand, the number of historical flood events in a given region is not large enough to build a robust statistical distribution of rainfall events and their associated flood depths. Also, depending on the rate of land use change, historical data may not be adequate to estimate flood risk under current conditions. On the other hand, generating a large number of simulated flood scenarios using H&H models is computationally expensive and would not be feasible. An alternative to historical data is represented by synthetic high-resolution flood scenarios generated using generative models.
[…]
@@
To address these challenges, we propose the Precipitation-Flood Depth Generative Pipeline, a novel methodology that harnesses generative machine learning to synthesize large-scale precipitation data and predict corresponding flood inundation depths through strategic sampling and modeling. The main workflow is illustrated in Fig. 1. Precipitation-Flood Depth Generative Pipeline starts with the training of a depth estimator using a number of H&H model-generated flood events, incorporating precipitation-based, spatial, and region-specific features. Following this, a Conditional GAN (CTGAN) with constraints is employed to generate synthetic rainfall precipitation events. Strategic thresholds are established to filter these synthetic records, ensuring close alignment with true precipitation patterns. For each location, synthetic events are smoothed using a K-nearest neighbors (KNN) algorithm and processed through the trained depth estimator to derive synthetic depth distributions. By iterating this procedure across multiple synthetic rainfall events, we construct flood probability maps for different combinations of rainfall characteristics (duration, peak precipitation intensity, and cumulative precipitation). These maps are validated using similarity and correlation metrics to confirm the correspondence of synthetic depth distributions to training data.
The novel contributions of the model presented in this paper are threefold:
@@
    We introduce a novel methodology that utilizes a surrogate machine learning pipeline to estimate flood probabilities using synthetic precipitation-flood events generated by CTGAN and depth estimator.
@@
Figure 1: The main research workflow comprises four key steps. Step I involves training and selecting the optimal depth estimation model to generate synthetic flood depth data. The trained model is later applied in Step III for synthetic flood depth construction. Step II focuses on sampling events, preprocessing point-level features, and training a CTGAN with specific constraints to generate a synthetic point cloud through conditional sampling. In Step III, a precipitation feature-level pool is created for each cell mesh using thresholds derived from sampled training event features, enabling strategic sampling to generate synthetic precipitation event features. The trained depth estimator from Step I then processes KNN-smoothed synthetic precipitation features along with static features to produce synthetic flood depths. Finally, Step IV iterates Step III to generate thousands of synthetic precipitation events, forming a synthetic depth distribution and ultimately constructing a synthetic flood probability map.
Flood-Precip GAN provides a scalable solution for generating high-resolution flood probability maps, to support to flood preparedness and mitigation efforts. The following section explains the components of the Flood-Precip GAN model and the associated datasets.
@@ 2.1.1 Study Region
Figure 2: Study area and Harris County Flood Control District gauge distribution. Map generated using ArcGIS Pro 3.0.0 (https://pro.arcgis.com/).
As shown in Fig. 2, Harris County, Texas, serves as the test bed for developing the Precipitation-Flood Depth Generative Pipeline. The primary watersheds in this area are Cypress Creek and Buffalo Bayou. Harris County, located at the heart of the Greater Houston Metropolitan Statistical Area, covers 1,778 square miles (4,605 square kilometers) and has seen its population grow to over 4.5 million in the past decade. The county’s flat topography, with elevations ranging from -40 feet (-12.19 meters) to 300 feet (91.44 meters), coupled with its largely developed land (over two-thirds), along with 20% pasture and cultivated lands 5, contributes to significant flood risk. The Cypress Creek watershed in the northern part and the Buffalo Bayou system in the central and southern areas flow into the San Jacinto River and the Ship Channel, eventually draining into the Gulf of Mexico. Dense urban development, poor natural drainage, limited soil infiltration, and the county’s subtropical climate further exacerbate its vulnerability to chronic flooding.
@@ 2.1.2 Data
In creating Precipitation-Flood Depth Generative Pipeline, there is a need for an initial set of rainfall events and related flood inundation scenarios. The lack of sufficient historical precipitation flood inundation data was addressed using 592 precipitation events processed with a physical-based model 9 to obtain corresponding flood scenarios. Specifically, 9 considered storm events occurred in Harris County in the period 2014-2022, and applied Rasterized Time-series Resampling Method (RTRM) to build 592 synthetic, realistic storms varying by spatial extent, duration, and rainfall intensity. 8 modeled those events in HEC-RAS 2D and produced as many flood scenarios. The study region in HEC-RAS 2D was modeled using a 1,200 ft $\times$ 1,200 ft mesh grid, with additional refinements applied near major rivers and tributaries to enhance both computational accuracy and stability. This final mesh consists of 26,301 individual cells. Manning’s roughness coefficient and imperviousness values were sourced from the 2019 National Land Cover Dataset 5, while soil infiltration values were obtained from the Gridded Soil Survey Geographic database 28. Model validation involved analyzing data from Harris County Flood Control District gauges 10 during five significant flood events from 2016 to 2020, including the 2016 Tax Day flood and Hurricane Harvey in 2017. This validation process, which calibrated the HEC-RAS 2D model for global minimum error across all gauges, confirmed the model’s reliability and accuracy for simulating diverse flood scenarios 8. We used the 592 rainfall events as the initial set from which to generate further synthetic rainfall events, as described in Section 2.4 and Section 2.5. The flood depth scenarios became the input for the depth estimator, a surrogate machine learning model for determining cell-wise maximum flood depth, described in Section 2.3.
@@ 2.2 Preprocessing
We employed Voronoi tessellation 14 to partition the training dataset, assigning each point within the study area a unique polygonal region to ensure there is no overlap among cells. Additionally, we segmented the 592 rainfall events used in the HEC-RAS 2D flood scenarios into three classes by inundation depth level, to ensure consistent stratification for subsequent dataset splitting and resampling. Specifically, for the training set, events are classified with mean depths of $\leq 2\ \text{inches}$, between $>2$ and $\leq 6\ \text{inches}$, and $>6\ \text{inches}$ in a ratio of 2:4:1, respectively. The validation set consists of 20% of the training set volume, with the remaining data designated as the test set. This approach optimizes the separation of events to ensure relatively balanced representation across the training, validation, and testing phases, facilitating more effective modeling.
@@ 2.3 Depth Estimator Model
For the depth estimator component of Precipitation-Flood Depth Generative Pipeline, we chose a tree-based machine learning model trained on simulated scenarios from the HEC-RAS 2D model. Compared to traditional machine learning models, tree-based models have shown superior performance due to their ability to manage complex, non-linear relationships and interactions within the data. XGBoost 3, 25, in particular, excels with its gradient boosting framework, which iteratively improves model accuracy by minimizing residual errors from previous iterations. Additionally, deep learning models in regression, such as the Transformer 29, 2, utilize attention mechanisms to capture long-range dependencies and complex feature interactions, providing substantial benefits in predictive performance. In this study, we employed the XGBoost regressor and the regression Transformer as universal depth estimators, training them on physics-based model-generated events that include both spatial features and precipitation-based features, as shown in Table 1. This integration aims to enhance the accuracy and robustness of maximum flood inundation estimations. Furthermore, we compared these universal models with a cell-wise model, MaxFloodCast V2 15, an XGBoost-based depth estimator aggregating 26,301 machine learning models, one for each of the 26,301 cells of the mesh. In MaxFloodCast V2 each cell model is trained independently of the others, without considering precipitation or water levels in neighboring cells. However, spatial dependencies exist in the natural water flow. For example, the water level of cells located in a channel is influenced by the water draining from upstream areas. To account for this spatial influence while maintaining cell-wise model independence, we introduced a heavy cumulative precipitation ratio (HCPR) and a heavy peak precipitation ratio (HPPR). Heavy precipitation is the hourly peak intensity or cumulative rain volume above a certain threshold established for a study area. HCPR and HPPR are defined at the watershed level as in Eq. 1 and Eq. 2.
$\text{R}_{i}^{c}=\frac{\sum_{j}AC_{i,j}\times h^{c}_{i,j}}{AW_{i}},c\in\{\mathfrak{c},\mathfrak{p}\}$ (1)
$h_{i,j}^{c}=\begin{cases}1,&\text{if }p_{i,j}^{c}>2\text{ in},\\ 0,&\text{otherwise},\end{cases}$ (2)
where $c$ represents HCPR as $\mathfrak{c}$ and HPPR as $\mathfrak{p}$, with $R$ denoting the respective heavy precipitation ratio. $AW_{i}$ refers to the total area of watershed $i$, while $AC_{i,j}$ represents the area of cell $j$ within watershed $i$. The binary variable $h_{i,j}$ identifies whether cell $j$ in watershed $i$ experiences heavy precipitation, and $p_{i,j}$ denotes the corresponding precipitation condition for cell $j$ within the same watershed. HCPR and HPPR are defined at the watershed scale, and are uniform across all cells within the same watershed during a rainfall event. These ratios can differ considerably across watersheds and across events, and offer substantial explanatory potential as surrogates for precipitation patterns.
@@
9 HCPRs (one for each watershed) Cell-wise Scalar (range 0 to 1) Proportion of watershed area experiencing heavy cumulative precipitation
9 HPPRs (one for each watershed) Cell-wise Scalar (range 0 to 1) Proportion of watershed area experiencing heavy peak precipitation
We assessed the predictive performance of the depth estimators (in Section 3.1) using Root Mean Squared Error (RMSE) and R-squared ($R^{2}$) scores 4 to evaluate both accuracy and model explainability. RMSE measures the average magnitude of the prediction errors, providing insight into the estimator’s precision, while $R^{2}$ indicates the proportion of variance in the observed data that is predictable from the input features, reflecting the model’s explanatory power. In addition, we conducted a comprehensive analysis of the aggregated performance of these models on both channel and non-channel cells. This distinction is crucial due to the varying hydrological and ground texture characteristics inherent to these environments, which influence the uncertainty and behavior of flood dynamics.